Ann, Ben, and Carol have some money. If Ann gives Ben 30 cents, then Ben will have twice as much as Ann. If Ben givens Carol 30 cents, then Carol will have as much as Ben. If Carol gives Ann 15 cents, they will both have the same amount of money. How much money to they each have?
Answer by richwmiller(17219)

2*(a-30)=(b+30)
b-30=(c+30)
c-15=a+15
a = 180, b = 270, c = 210